Questions 1 and 2 refer to the following information: X Company is considering buying a part next year that it currently makes. This years production costs for 3,200 units were: Direct materials Direct labor Variable overhead Fixed overhead Total Per-Unit 3.85 3.65 3.50 4.00 Total $12,320 11,680 11,200 12,800 $48,000 $15.00 A company has offered to supply this part for $14.85 per unit. 6,400 of X Companys fixed overhead are allocated costs that will occur even if they buy the part. But if X Company buys the part, it can rent out the freed-up resources for $2,100. Production next year is expected to be 3,700 units. 1. If X Company continues to make the part instead of buying it, it will save Submit Answer Tries 0/3 2. At what production level would X Company be indifferent between making and buying the part? Submit Answer Tries 0/3
Answer #1

1) Differential analysis

Make Buy
Direct material (3700*3.85) 14245
Direct labour (3700*3.65) 13505
Variable overhead (3700*3.50) 12950
Fixed overhead (12800-6400) 6400
Opportunity Cost 2100
Purchase cost (3700*14.85) 54945
Total relevant cost 49200 54945

it will save = 54945-49200 = $5745

2) Indifferent point

Make Cost = Purchase cost

11X+8500 = 14.85X

-3.85X = -8500

X(indifferent point) = 8500/3.85 = 2208 Units